OpenStudy (anonymous):

Find the first derivative for g(x) = 3(6 − 9x)^4 A. g′(x) = 108(6 − 9x)^3 B. g′(x) = −12(6 − 9x)^3 C. None of these D. g′(x) = −108(6 − 9x)^3 E. g′(x) = 12(6 − 9x)^3

OpenStudy (anonymous):

Its D. g′(x) = −108(6 − 9x)^3

OpenStudy (anonymous):

\[g(x) = 3(6 − 9x)^4 \rightarrow g'(x) = 3 \times 4 (6 − 9x)^3 \frac{d}{dx}(6 − 9x)\] \[= 12 (6 − 9x)^3 (-9) = 12(-9) (6 − 9x)^3 = -108 (6 − 9x)^3 \] Which is the required derivative. @lesliemariahp

OpenStudy (anonymous):

Thank you so much! @dpasingh